### v1

@@ 3.3 Design of Validation Objective
In this section, we take into consideration two principles for the construction of validation objectives. First, the validation loss should mimic the behaviour of the model on the testing data stream. Assuming the limited labeled data takes a snapshot of the whole testing data stream and classification is the task to be addressed, an intuitive choice is the cross-entropy loss as follows.
$\mathcal{H}_{ce}(x_{i};\theta^{*}_{m})=\sum_{c=1\cdots C}\mathbbm{1}(y_{i}=c)\log h_{c}(x_{i};\theta^{*}_{m})$
@@
Regularization for Stable Model Selection: The stability of model selection heavily depends on the selection of validation set and the variation of target domain distribution. Under a more realistic TTA scenario, the target domain distribution could shift over time, e.g. the continual test-time adaptation [Wang et al.(2022)Wang, Fink, Van Gool, and Dai]. Such a realistic challenge may render cross-entropy loss less effective due to over-adapting to a local distribution. To stabilize the model selection procedure, we propose the following two measures.
@@
Anchor Deviation for Regularizing Model Selection: We first introduce an anchor deviation to regularize the model selection procedure. This regularization is similar to the anchor network proposed in  [Su et al.(2024a)Su, Xu, and Jia] except that we use the anchor deviation to regularize model selection which is orthogonal to the purpose in [Su et al.(2024a)Su, Xu, and Jia]. Specifically, we keep a frozen source domain model, denoted as $\theta_{0}$. The anchor loss is defined as the L2 distance between the posteriors of the frozen source model and the $m$-th candidate model $\theta^{*}_{m}$. A big deviation from the source model prediction suggests the model has been adapted to a very specific target domain and further adapting towards one direction would increase the risk of failing to adapt to the continually changing distribution.
$\mathcal{R}_{anc}(x_{i};\theta^{*}_{m})=\frac{1}{N_{b}}\sum\limits_{x_{i},y_{i}\in{D}^{l}}||h(x_{i};\theta_{0})-h(x_{i};\theta^{*}_{m})||$
@@
Smoothed Scoring for Model Selection: To determine the optimal model to select, we combine the cross-entropy loss and the anchor deviation as the final score. Direct summing the two metrics as the model selection score is sub-optimal due to the relative scale between the two metrics. Therefore, we propose to first normalize the two metrics. After the normalization, the relative size relationships of different candidate model scores are preserved, and a comparable validation loss can be conveniently obtained by addition, as below.
@@
$\begin{split}&\forall m=1\cdots|\Omega|,\quad\;\mathrm{Norm}(x)=({x-\mathrm{min}(x)})/({\mathrm{max}(x)-\mathrm{min}(x)}),\quad S_{ce}(x_{i};\theta_{m}^{*})=\mathrm{Norm}(\mathcal{H}_{ce}(x_{i};\theta^{*}_{m}))\\ &S_{anc}(x_{i};\theta^{*}_{m})=\mathrm{Norm}(\mathcal{R}_{anc}(x_{i};\theta^{*}_{m})),\quad\mathcal{L}_{val}(x_{i};\theta^{*}_{m})=S_{ce}(x_{i};\theta^{*}_{m})+S_{anc}(x_{i};\theta^{*}_{m})\end{split}$
@@
Considering that testing data stream is often highly correlated, i.e. temporally adjacent samples are likely subject to the same distribution shift, the optimal hyper-parameter/model should change smoothly as well. To encourage a smooth transition of the optimal model, we select the optimal model following an exponential moving average fashion as in Eq. 6 where we denote the moving average validation loss at the $t$-th batch as ¯$\bar{\mathcal{L}}_{val}^{t}$.
¯$\begin{split}&\bar{\mathcal{L}}_{val}^{t}(\theta_{m}^{*})=\beta\bar{\mathcal{L}}_{val}^{t-1}(\theta_{m}^{*})+(1-\beta)\frac{1}{|{B}_{t}^{l}|}\sum_{x_{j},y_{j}\in{B}_{t}^{l}}\mathcal{L}_{val}(h(x_{j};\theta_{m}^{*}),y_{j}),\quad\theta^{*}=\arg\min_{m}\bar{\mathcal{L}}_{val}^{t}(\theta_{m}^{*})\end{split}$
@@ 3.4 Sample Selection for HILTTA
The sample selection strategy should balance the objectives of active learning and model selection for more effective HILTTA. We achieve this target from two perspectives. First, the selected samples should prioritize low-confidence samples for more effective active learning, as adopted by SimATTA [Gui et al.(2024)Gui, Li, and Ji] through incremental K-means. However, selecting redundant low-confidence samples may harm the effectiveness of model validation. This motivates us to select samples approximating the testing data distribution, similar to the objective of active testing [Kossen et al.(2021)Kossen, Farquhar, Gal, and Rainforth]. Biasing towards either criterion would result in suboptimal HILTTA.
The above objectives could be achieved by a recent active learning approach, BADGE [Ash et al.(2020)Ash, Zhang, Krishnamurthy, Langford, and Agarwal], which takes the gradient of loss w.r.t. the penultimate layer feature as the uncertainty weighted feature as $g_{x}=\frac{\partial}{\partial\theta_{\theta_{out}}}\mathcal{L}_{ce}(h(x;\theta),\hat{y})$, on which K-Means++ is applied for active selection. The feature dimension is the product of the number of classes $C$ and penultimate layer channels $D$, i.e. $g_{x}\in\mathbbm{R}^{C\times D}$. Accumulating the feature over all categories results in overly high dimension features for large-scale classification tasks, e.g. $C=1000$ for ImageNet [Deng et al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei]. To mitigate the ultra-high dimension challenge, we integrate uncertainty into feature representation by multiplying the uncertainty, measured by the margin confidence, to the feature representation as in Eq. 7, where $\mathrm{sort}^{d}$ refers to the descending sorting.
$\begin{split}&\hat{p}=\mathrm{sort}^{d}[h_{1}(x_{i};\theta),h_{2}(x_{i};\theta),\cdots h_{C}(x_{i};\theta)],&g_{i}=(1-\hat{p}_{1}+\hat{p}_{2})f(x_{i};\theta)\end{split}$
@@
Sample Selection via K-Margin: The samples to be selected for annotation are eventually obtained by conducting a K-Center clustering on the uncertainty-weighted features within each minibatch, as in Eq. 8, which is referred to as K-Margin clustering. We select K samples ${B}_{t}^{l}$ such that the maximal distance of all testing sample embedding is minimized. We deliberately exclude the selected samples in the previous minibatch, as opposed to SimATTA [Gui et al.(2024)Gui, Li, and Ji], because the labeled samples will serve as the validation set, and excluding new samples that may overlap with previously selected samples could undermine the effectiveness of model selection. The following problem can be solved via a greedy selection algorithm [Sener and Savarese(2018)].
$\min_{{B}_{t}^{l}\subseteq\{g_{i}\}_{i=1}^{N_{b}}}\max_{g_{i}\in\{g_{i}\}_{i=1}^{N_{b}}}\min_{c_{k}\in{B}_{t}^{l}}||g_{i}-c_{k}||,\quad s.t.\;|{B}_{t}^{l}|\leq K$ (8)
@@ 4.1 Experimental setting
Datasets: We select four datasets for evaluation. The CIFAR10-C and CIFAR100-C [Hendrycks and Dietterich(2018)] are small-scale corruption datasets, with 15 different common corruptions, each containing 10,000 corrupt images with 10/100 categories. For our evaluation in large-scale datasets, we opt for ImageNet-C [Hendrycks and Dietterich(2018)], which also contains 15 different corruptions, each with 50,000 corrupt images in 1000 categories. Additionally, ImageNet-D [Rusak et al.(2022)Rusak, Schneider, Gehler, Bringmann, Brendel, and Bethge] is a style-transfer dataset, offering 6 domain shifts, each consisting of 10,000 images selected in 109 classes.
Implementation Details: We follow the continual TTA setting [Wang et al.(2022)Wang, Fink, Van Gool, and Dai, Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan] to evaluate TTA performance under long-term adaptation, where the target domain undergoes continuous changes. For CIFAR10-C and CIFAR100-C, we use a batch size of 200 and the annotation percentage of $3\%$, with respectively employed pre-trained WideResNet-28 [Zagoruyko and Komodakis(2016)] and ResNeXt-29 [Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] models. For ImageNet-C and ImageNet-D, we use a batch size of 64 for adaptation, and the annotation percentage is 3.2% of the total testing samples, with the ResNet-50 [He et al.(2016)He, Zhang, Ren, and Sun] pretrained model. For unsupervised training, we adopt the optimizer for the respective methods presented in the original paper. For supervised training, we use Adam optimizer, with a 1e-5 learning rate for all methods throughout all datasets. We consider seven distinct values for each hyper-parameter category regarding model selection. We defer more detailed experimental settings to the appendix.
@@ 4.2 Evaluations on HILTTA
Evaluation of Active Learning & Model Selection Strategies: We report the classification error on the four datasets in Tab. 1. Each method is specified by two attributes. The “Active Learning” attribute indicates the sample selection method for additional human annotation. The “Model Selection” attribute indicates whether the method enables hyper-parameter selection proposed in Sec. 3.2. We use TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] as an off-the-shelf TTA method. For methods that do not enable ”Model Selection”, we report three results, the worst (Worst), average (Avg.), and best (Best). The worst or best results are obtained by a fixed hyper-parameter trained with which the model exhibits the worst or best overall accuracy. The average results are obtained by averaging the overall accuracy across the models trained upon all candidate hyper-parameters. We make the following observations from the results in Tab. 1. i) Without additional annotation (TENT) the worst and average performance could vary substantially (best 62.96% v.s. worst 95.23% on ImageNet-C). The results suggest the sensitivity to hyper-parameter remains a major challenge. ii) With additional annotation (TENT w/ HIL), we first notice that applying active learning alone consistently improves over fully unsupervised adaptation in terms of worst, average and best results. iii) When model selection is further applied, the hyper-parameter is automatically chosen, resulting in a single error rate. We witness that for all active learning methods, our model selection strategy achieves significant improvement from the average error rate. In many cases, we could even surpass the best result. iv) Finally, our proposed K-Margin selection strategy combined with model selection almost outperforms all other active learning techniques, suggesting the effectiveness of balancing active learning and model selection for HILTTA.
Active Model ImageNet-C ImageNet-D CIFAR100-C CIFAR10-C
Learning Select. Worst Avg. Best Worst Avg. Best Worst Avg. Best Worst Avg. Best
Source $\times$ $\times$ 82.02 58.98 46.44 43.51
TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] $\times$ $\times$ 95.23 73.13 62.96 83.54 59.82 52.93 94.26 50.69 32.75 71.22 27.77 18.14
$\times$ 95.04 67.08 (+6.05) 59.60 79.15 55.01 (+4.81) 49.59 94.26 50.70 (-0.01) 30.87 64.61 24.91 (+2.86) 16.84
Random 59.42 (+13.71) 49.56 (+10.26) 31.50 (+19.19) 17.80 (+9.97)
$\times$ 95.23 70.47 (+2.66) 60.79 82.39 56.27 (+3.55) 49.85 94.37 51.33 (-0.64) 30.18 29.46 18.45 (+9.32) 15.77
Entropy [Shannon(1948)] 60.45 (+12.68) 49.48 (+10.34) 32.78 (+17.91) 15.82 (+11.95)
$\times$ 94.49 69.90 (+3.23) 60.73 80.15 56.40 (+3.42) 50.83 94.56 53.08 (-2.39) 32.08 60.62 25.92 (+1.85) 17.81
K-Center [Sener and Savarese(2018)] 60.56 (+12.57) 50.60 (+9.22) 33.25 (+17.44) 17.97 (+9.80)
TENT $\times$ 93.82 71.51 (+1.62) 62.78 79.06 54.91 (+4.91) 49.65 94.34 51.20 (-0.51) 31.59 59.06 24.41 (+3.36) 17.15
w/ HIL VeSSAL [Saran et al.(2023)Saran, Yousefi, Krishnamurthy, Langford, and Ash] 62.37 (+10.76) 49.56 (+10.26) 32.89 (+17.80) 17.81 (+9.96)
$\times$ 93.57 68.85 (+4.28) 60.66 78.32 54.25 (+5.57) 49.35 94.35 48.19 (+2.50) 30.53 70.23 25.36 (+2.41) 17.04
ASE [Kossen et al.(2022)Kossen, Farquhar, Gal, and Rainforth] 60.52 (+12.61) 53.08 (+6.74) 31.16 (+17.53) 17.49 (+10.28)
$\times$ 94.59 69.09 (+4.04) 60.71 79.99 55.25 (+4.57) 49.89 94.64 48.70 (+1.99) 30.36 31.28 19.35 (+8.42) 16.39
SimATTA [Gui et al.(2024)Gui, Li, and Ji] 60.46 (+12.67) 49.59 (+10.23) 31.23 (+19.46) 16.95 (+10.82)
K-Margin $\times$ 94.98 68.13 (+5.00) 58.49 76.27 53.77 (+6.05) 48.55 94.75 49.37 (+1.32) 30.11 27.81 18.21 (+9.56) 15.75
(OURS) 58.35 (+14.78) 48.74 (+11.08) 30.53 (+20.16) 15.87 (+11.90)
@@
Table 1: Evaluation under HILTTA with average classification error (lower is better). ’Worst’, ’Avg.’, and ’Best’ denote the worst, average, and best performances, respectively, with the candidate hyper-parameter pool $\Omega$. The performance improvement by Active Learning and Model Selection is denoted within brackets as ($\delta$).
Evaluation under Alternative TTA Methods: Moreover, we investigate various off-the-shelf TTA methods in Tab. 2. Employing our proposed K-Margin as the sample selection method, we consistently observe enhancements for all off-the-shelf TTA methods when paired with our HIL model selection and learning strategies. These findings advocate for our active learning and model selection approach regardless of TTA methods. Importantly, even the simplest baseline, TENT, achieves comparable results to state-of-the-art TTA methods under HILTTA. The results suggest exploiting small amount of labeled data is very effective for TTA and the margin between different TTA methods are smaller than expected when equipped with human-in-the-loop.
TTA A.L. M.S. ImageNet-C ImageNet-D CIFAR100-C CIFAR10-C
Methods Worst Avg. Best Worst Avg. Best Worst Avg. Best Worst Avg. Best
$\times$ $\times$ 95.23 73.13 62.96 83.54 59.82 52.93 94.26 50.69 32.75 71.22 27.77 18.14
TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] $\times$ 94.98 68.13 (+5.00) 58.49 76.27 53.77 (+6.05) 48.55 94.75 49.37 (+1.32) 30.11 27.81 18.21 (+9.56) 15.75
58.35 (+14.78) 48.74 (+11.08) 30.53 (+20.16) 15.87 (+11.90)
$\times$ $\times$ 89.78 81.48 67.75 63.06 56.59 52.45 37.31 34.67 32.76 19.95 19.11 18.54
PL [Lee et al.(2013)] $\times$ 79.14 70.30 (+11.18) 62.15 50.93 49.25 (+7.34) 48.15 32.08 30.85 (+3.82) 29.93 17.02 16.67 (+2.44) 16.45
61.60 (+19.88) 49.01 (+7.58) 30.82 (+3.85) 16.97 (+2.14)
$\times$ $\times$ 99.18 79.56 64.68 91.51 64.51 53.69 90.72 43.89 32.90 52.87 23.22 17.26
SHOT [Liang et al.(2020)Liang, Hu, and Feng] $\times$ 99.29 74.49 (+5.07) 59.30 90.86 58.19 (+6.32) 48.37 87.30 41.05 (+2.84) 30.51 21.27 16.68 (+6.54) 15.66
60.66 (+18.90) 48.98 (+15.53) 31.13 (+12.76) 15.78 (+7.44)
$\times$ $\times$ 99.81 70.85 57.92 94.68 60.89 52.00 44.89 34.57 31.32 39.28 21.66 17.14
EATA [Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan] $\times$ 99.66 68.36 (+2.49) 54.37 94.75 56.93 (+3.96) 47.88 42.10 32.42 (+2.15) 29.51 25.06 17.82 (+3.84) 15.69
56.75 (+14.10) 49.17 (+11.72) 30.43 (+4.14) 15.79 (+5.87)
$\times$ $\times$ 90.60 70.70 62.20 56.19 53.37 51.72 39.00 33.84 31.90 20.22 19.10 17.45
SAR [Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan] $\times$ 89.03 64.81 (+5.89) 58.22 49.79 48.64 (+4.73) 47.50 32.24 30.63 (+3.21) 29.67 17.14 16.60 (+2.50) 15.63
58.91 (+11.79) 50.02 (+3.35) 30.00 (+3.84) 16.87 (+2.23)
$\times$ $\times$ 98.77 72.28 60.45 88.80 58.92 50.14 38.35 32.18 30.16 29.36 19.75 16.66
RMT [Döbler et al.(2023)Döbler, Marsden, and Yang] $\times$ 98.50 69.87 (+2.41) 56.67 87.96 55.51 (+3.41) 46.12 37.68 31.26 (+0.92) 29.02 27.04 18.35 (+1.40) 15.80
58.26 (+14.02) 46.28 (+12.64) 29.41 (+2.77) 16.45 (+3.30)
Table 2: Evaluation of different off-the-shelf TTA methods with K-Margin selection method under HILTTA protocol.

### v2

@@ 3.3 Design of Validation Objective
In this section, we take into consideration two principles for the construction of validation objectives. First, the validation loss should mimic the behavior of the model on the testing data stream. Assuming the limited labeled data takes a snapshot of the whole testing data stream and classification is the task to be addressed, an intuitive choice is the cross-entropy loss as follows.
$\mathcal{H}_{ce}(x_{i};\theta^{*}_{m})=\sum_{c=1}^{C}\mathbbm{1}(y_{i}=c)\log h_{c}(x_{i};\theta^{*}_{m})$ (3)
@@
Regularization for Stable Model Selection: The stability of model selection heavily depends on the selection of the validation set and the variation of target domain distribution. Under a more realistic TTA scenario, the target domain distribution could shift over time, e.g. the continual test-time adaptation (Wang et al., 2022). Such a realistic challenge may render cross-entropy loss less effective due to over-adapting to local distribution. To stabilize the model selection procedure, we propose the following two measures.
@@
Anchor Deviation for Regularizing Model Selection: We first introduce an anchor deviation to regularize the model selection procedure. This regularization is similar to the anchor network proposed in  Su et al. (2024a) except that we use the anchor deviation to regularize model selection which is orthogonal to the purpose in Su et al. (2024a). Specifically, we keep a frozen source domain model, denoted as $\theta_{0}$. The anchor loss is defined as the L2 distance between the posteriors of the frozen source model and the $m$-th candidate model $\theta^{*}_{m}$. A big deviation from the source model prediction suggests the model has been adapted to a very specific target domain and further adapting towards one direction would increase the risk of failing to adapt to the continually changing distribution.
$\mathcal{R}_{anc}(x_{i};\theta^{*}_{m})=\frac{1}{N_{b}}\sum_{(x_{i},y_{i})\in\mathbb{D}^{l}}\left\|h(x_{i};\theta_{0})-h(x_{i};\theta^{*}_{m})\right\|$ (4)
@@
Smoothed Scoring for Model Selection: To determine the optimal model to select, we combine the cross-entropy loss and the anchor deviation as the final score. Direct summing the two metrics as the model selection score is sub-optimal due to the relative scale between the two metrics. Therefore, we propose to first normalize the two metrics as $\mathrm{Norm}(x)=\frac{x-\mathrm{min}(x)}{\mathrm{max}(x)-\mathrm{min}(x)}$. After the normalization, the relative size relationships of different candidate model scores are preserved, and a comparable validation loss can be conveniently obtained by addition, as below.
@@
$\begin{split}&\mathcal{L}_{val}(x_{i};\theta_{m}^{*})=S_{ce}(x_{i};\theta_{m}^{*})+S_{anc}(x_{i};\theta_{m}^{*})\\ \text{s.t.}\quad&\forall m=1,\ldots,|\Omega|,\quad S_{ce}(x_{i};\theta_{m}^{*})=\mathrm{Norm}(\mathcal{H}_{ce}(x_{i};\theta^{*}_{m})),~{}~{}S_{anc}(x_{i};\theta_{m}^{*})=\mathrm{Norm}(\mathcal{R}_{anc}(x_{i};\theta^{*}_{m})),\\ \end{split}$ (5)
@@
Considering that the testing data stream is often highly correlated, i.e. temporally adjacent samples are likely subject to the same distribution shift, the optimal hyper-parameter/model should change smoothly as well. To encourage a smooth transition of the optimal model, we select the optimal model following an exponential moving average fashion as in Eq. 6 where we denote the moving average validation loss at the $t$-th batch as ¯$\bar{\mathcal{L}}_{val}^{t}$ and the chosen best model $\theta^{*}$ is obtained as the one that minimizes the smoothed validation loss.
¯$\begin{split}\bar{\mathcal{L}}_{val}^{t}(\theta_{m}^{*})&=\beta\bar{\mathcal{L}}_{val}^{t-1}(\theta_{m}^{*})+(1-\beta)\frac{1}{|\mathbb{B}_{t}^{l}|}\sum_{x_{j},y_{j}\in\mathbb{B}_{t}^{l}}\mathcal{L}_{val}(h(x_{j};\theta_{m}^{*}),y_{j}),\\ \theta^{*}&=\arg\min_{m}\bar{\mathcal{L}}_{val}^{t}(\theta_{m}^{*})\end{split}$ (6)
@@ 3.4 Sample Selection for HILTTA
The sample selection strategy should balance the objectives of active learning and model selection for more effective HILTTA. We achieve this target from two perspectives. First, the selected samples should prioritize low-confidence samples for more effective active learning, as adopted by SimATTA (Gui et al., 2024) through incremental K-means. However, selecting redundant low-confidence samples may harm the effectiveness of model validation. This motivates us to select samples approximating the testing data distribution, similar to the objective of active testing (Kossen et al., 2021). Biasing towards either criterion would result in a suboptimal HILTTA.
The above objectives could be achieved by a recent active learning approach, BADGE (Ash et al., 2020), which takes the gradient of loss w.r.t. the penultimate layer feature as the uncertainty weighted feature as $g_{x}=\frac{\partial}{\partial\theta_{\theta_{out}}}\mathcal{L}_{ce}(h(x;\theta),\hat{y})$, on which K-Means++ is applied for active selection. The feature dimension is the product of the number of classes $C$ and penultimate layer channels $D$, i.e. $g_{x}\in\mathbbm{R}^{C\times D}$. Accumulating the feature overall categories results in overly high dimension features for large-scale classification tasks, e.g. $C=1000$ for ImageNet (Deng et al., 2009). To mitigate the ultra-high dimension challenge, we integrate uncertainty into feature representation by multiplying the uncertainty, measured by the margin confidence, to the feature representation as in Eq. 7, where $\mathrm{sort}^{d}$ refers to the descending sorting.
$\begin{split}\hat{p}&=\text{sort}^{d}\left[h_{1}(x_{i};\theta),h_{2}(x_{i};\theta),\cdots,h_{C}(x_{i};\theta)\right],\\ g_{i}&=(1-\hat{p}_{1}+\hat{p}_{2})f(x_{i};\theta)\end{split}$ (7)
@@
Sample Selection via K-Margin: The samples to be selected for annotation are eventually obtained by conducting a K-Center clustering on the uncertainty-weighted features within each minibatch, as in Eq. 8, which is referred to as K-Margin clustering. We select K samples $\mathbb{B}_{t}^{l}$ such that the maximal distance of all testing sample embedding is minimized. We deliberately exclude the selected samples in the previous minibatch, as opposed to SimATTA (Gui et al., 2024), because the labeled samples will serve as the validation set, and excluding new samples that may overlap with previously selected samples could undermine the effectiveness of model selection. The following problem can be solved via a greedy selection algorithm (Sener & Savarese, 2018).
$\min_{\mathbb{B}_{t}^{l}\subseteq\{g_{i}\}_{i=1}^{N_{b}}}\max_{g_{i}\in\{g_{i}\}_{i=1}^{N_{b}}}\min_{c_{k}\in\mathbb{B}_{t}^{l}}||g_{i}-c_{k}||,\quad s.t.\;|\mathbb{B}_{t}^{l}|\leq K$ (8)
3.5 Overall Algorithm for HILTTA
We propose two types of training losses for model updates and a validation loss for model selection. During the model selection stage, we generate multiple candidate models by training with an unsupervised TTA loss, denoted as $\mathcal{L}_{tr}^{u}$, following the off-the-shelf TTA methods. The optimal candidate model $\theta^{t*}$ is selected using $\arg\min\limits_{m}\bar{\mathcal{L}}_{val}^{t}(\theta^{*}_{m})$. After selecting the best model, we further refine it through supervised training, optimizing a cross-entropy loss $\mathcal{L}_{tr}^{l}$. The complete Human-in-the-Loop Test-Time Adaptation algorithm is detailed in Alg. 1.
Input : Source Model $\theta_{0}$; Candidate Hyper-parameters $\Omega=\{\omega_{m}\}$; Testing Data Batches $\{\mathbb{B}_{t}\}$
Output : Predictions $\hat{\mathcal{Y}}=\{\hat{y}_{i}\}$
for $t=1$ to $T$ do
       # Make Predictions:
       $\forall x_{i}\in\mathbb{B}_{t},\;\hat{y}_{i}=h(x_{i};\theta^{t-1*}),\;\hat{\mathcal{Y}}=\hat{\mathcal{Y}}\cup\hat{y}_{i}$
       # Oracle Annotation:
       Select labeled subset $\mathbb{B}_{t}^{l}$ by Eq. 8
       for $m=1$ to $N_{m}$ do
             # Unsupervised Model Adaptation:
             $\theta_{m}^{t*}=\arg\min\limits_{\theta}\frac{1}{N_{b}}\sum\limits_{x_{i}\in\mathbb{B}_{t}^{u}}\mathcal{L}_{tr}^{u}(h(x_{i};\theta);\omega_{m})$
      # Update Moving Average Validation Loss:
       Update ¯$\bar{\mathcal{L}}_{val}^{t}$ by Eq. 6.
       # Model Selection by Eq. 6:
       $\theta^{t*}=\arg\min\limits_{m}\bar{\mathcal{L}}_{val}^{t}(\theta^{*}_{m})$
       # Supervised Model Adaptation:
       $\theta^{t*}=\arg\min_{\theta}\frac{1}{N_{b}}\sum\limits_{x_{i},y_{i}\in\mathbb{B}_{t}^{l}}\mathcal{L}_{tr}^{l}(h(x_{i};\theta),y_{i};\omega_{m})$
return Predictions $\hat{\mathcal{Y}}$;
Algorithm 1 Human-in-the-Loop TTA
@@ 4.1 Experimental setting
Datasets: We select a total of five datasets for evaluation. The CIFAR10-C and CIFAR100-C (Hendrycks & Dietterich, 2018) are small-scale corruption datasets, with 15 different common corruptions, each containing 10,000 corrupt images with 10/100 categories. For our evaluation in large-scale datasets, we opt for ImageNet-C (Hendrycks & Dietterich, 2018), which also contains 15 different corruptions, each with 50,000 corrupt images in 1000 categories. Additionally, ImageNet-D (Rusak et al., 2022) is a style-transfer dataset, offering 6 domain shifts, each consisting of 10,000 images selected in 109 classes. Additionally, we evaluate our method on the ModelNet40-C dataset (Sun et al., 2022), which includes 3,180 3D point clouds affected by 15 common types of corruption.
Implementation Details: We evaluate TTA performance in the continual adaptation setting (Wang et al., 2022; Niu et al., 2022), where the target domain undergoes continuous changes. For CIFAR-10-C and CIFAR100-C datasets, we use a batch size of 200 and an annotation percentage of 3%, with pre-trained WideResNet-28 (Zagoruyko & Komodakis, 2016) and ResNeXt-29 (Xie et al., 2017) models, respectively. For ImageNet-C and ImageNet-D, we use a batch size of 64 for adaptation and an annotation percentage of 3.2% of the total testing samples, employing the ResNet-50 (He et al., 2016) pre-trained model.
We use the optimizer recommended in the original papers for each method’s unsupervised training. For supervised training, we use the Adam optimizer with a learning rate of 1e-5. We consider seven distinct values for each hyper-parameter category regarding model selection, detailed in Tab. 6.
Competing Methods: We evaluate 6 state-of-the-art off-the-shelf test-time adaptation approaches under both traditional unsupervised TTA and Human-in-the-Loop TTA protocols, including TENT (Wang et al., 2020), PL (Lee et al., 2013), SHOT (Liang et al., 2020), EATA (Niu et al., 2022), SAR (Niu et al., 2023), RMT (Döbler et al., 2023). For each of the above TTA method under HILTTA protocol, we embed the TTA loss into the “Unsupervised Model Adaptation” step of the HILTTA algorithm in Alg. 1. We further benchmark against existing active TTA methods, including SimATTA (Gui et al., 2024) and VeSSAL (Saran et al., 2023). Specifically, we adapt VeSSAL for TTA by using the selected samples for supervised training. More details of the benchmarked methods are deferred to the Appendix.
@@ 4.2 Evaluations on HILTTA
We report the classification error rates for continual TTA in Tab.1. For unsupervised TTA methods (Unsup. TTA), we report three results: worst(Worst), average (Avg.), and best (Best). The "Worst" and "Best" results are derived from fixed hyper-parameters that yield the highest and lowest overall error rate, respectively. The "Average" results are calculated by averaging the accuracy across models trained with all candidate hyper-parameters. For the competing active TTA methods (Active TTA), we maintain the same annotation rate as our method and use manually tuned hyper-parameters to ensure a fair comparison. These methods, lacking model selection capabilities, result in a single error rate per dataset. In contrast, our method (Human-in-the-Loop TTA) automatically selects hyper-parameters using our proposed model selection strategy, also resulting in a single error rate per dataset.
@@
Table 1: Average classification error (lower is better) through ongoing adaptation to varying domains. "Worst", "Avg.", and "Best" indicate the worst, average, and best performance within the candidate hyper-parameter pool $\Omega$. The improvement of error rate for each off-the-shelf TTA method after combining with HIL is appended in the bracket as ($\delta$). * denotes methods under manually tuned hyper-parameters. $\diamondsuit$ denotes methods combined with our proposed HILTTA (with model selection).
TTA Method ImageNet-C (%) $\downarrow$ ImageNet-D (%) $\downarrow$ CIFAR100-C (%) $\downarrow$ CIFAR10-C (%) $\downarrow$
Worst  Avg. Best Worst  Avg. Best Worst  Avg. Best Worst  Avg. Best
Unsup. TTA Source 82.02 58.98 46.44 43.51
TENT (Wang et al., 2020) 95.23 73.13 62.96 83.54 59.82 52.93 94.26 50.69 32.75 71.22 27.77 18.14
PL(Lee et al., 2013) 89.78 81.48 67.75 63.06 56.59 52.45 37.31 34.67 32.76 19.95 19.11 18.54
SHOT(Liang et al., 2020) 99.18 79.56 64.68 91.51 64.51 53.69 90.72 43.89 32.90 52.87 23.22 17.26
EATA(Niu et al., 2022) 62.41 60.26 58.26 51.75 51.18 50.78 51.30 42.72 32.34 20.17 19.57 17.58
SAR(Niu et al., 2023) 90.60 70.70 62.20 56.19 53.37 51.72 39.00 33.84 31.90 20.22 19.10 17.45
RMT(Döbler et al., 2023) 98.77 72.28 60.45 88.80 58.92 50.14 38.35 32.18 30.16 29.36 19.75 16.66
Active TTA* VeSSAL (Saran et al., 2023) 62.78 49.77 45.97 17.47
SimATTA (Gui et al., 2024) 64.09 49.26 31.75 17.66
Human-in- the-Loop TTA (OURS)$\diamondsuit$ HIL + TENT 58.35 (+14.78) 48.74 (+11.08) 30.53 (+20.16) 15.87 (+11.90)
HIL + PL 61.60 (+19.88) 49.01 (+7.58) 30.82 (+3.85) 16.97 (+2.14)
HIL + SHOT 60.66 (+18.90) 48.98 (+15.53) 31.13 (+12.76) 15.78 (+7.44)
HIL + EATA 55.13 (+5.13) 47.51 (+3.67) 30.93 (+11.79) 16.40 (+3.17)
HIL + SAR 58.91 (+11.79) 50.02 (+3.35) 30.00 (+3.84) 16.87 (+2.23)
HIL + RMT 58.26 (+14.02) 46.28 (+12.64) 29.41 (+2.77) 16.45 (+3.30)
From the results, we make the following observations:
i) Existing unsupervised TTA methods, when applied without additional annotation, exhibit significant sensitivity to hyper-parameters. This sensitivity can lead to large variations in performance (e.g., TENT’s best accuracy at 62.96% versus its worst at 95.23% on ImageNet-C). In the worst case, performance can even fall below the source model without any adaptation. In contrast, the "Best" performance across all methods are relatively close to each other. Nevertheless, selecting the best model without oracle knowledge is impractical in real-world scenarios. These findings highlight the persistent challenge of hyper-parameter sensitivity in TTA.
ii) Our proposed human-in-the-loop TTA consistently improves the performance of all off-the-shelf TTA methods against their unsupervised counterparts even with the best hyper-parameters. This demonstrates the robustness of our approach, mitigating the catastrophic failures that can arise from hyper-parameter sensitivity.
iii) Compared to existing active TTA approaches, even when manually tuning hyper-parameters to address the model selection issue, our proposed HILTTA consistently outperforms them across all four datasets. Notably, even a basic baseline like TENT achieves performance comparable to or better than existing active TTA methods when combined with our approach. These results highlight the effectiveness of integrating a small amount of labeled data through synergistic active learning and model selection, rather than relying solely on active learning.